OpenStudy (anonymous):

Find an equation for the nth term of the arithmetic sequence. 8, 6, 4, 2, ...

OpenStudy (jim766):

An=a1+ (n-1)d Do you know this formula?

OpenStudy (anonymous):

@jim766 yes but I don't understand it

OpenStudy (jim766):

A1 is the first term D is the common difference Common difference is the number you add each time to find the next number

OpenStudy (jim766):

So what is the common difference?

OpenStudy (anonymous):

@jim766 -2

OpenStudy (jim766):

Right, What is the first term?

OpenStudy (jim766):

What is the first number in the sequence?

OpenStudy (anonymous):

@jim766 8

OpenStudy (jim766):

An = 8 + ( n - 1) (-2)

OpenStudy (jim766):

Now distribute the -2

OpenStudy (anonymous):

@jim766 An = 8+ -2n + 2 = 10 + -2n

OpenStudy (jim766):

Right, nice job!